Mrs. De Leon is three times as old as her daughter. In twelve years, she will be only twice as old. How old are they now?
Answer by Cromlix(4381)

Mrs. De Leon is three times as old as her daughter.
MDL = 3D
In 12 years she will be only twice as old.
MDL + 12 = 2(D + 12)
Replace MDL with 3D from first equation.
3D + 12 + 2(D + 12)
3D + 12 = 2D + 24
Collect like terms
3D - 2D = 24 - 12
D = 12
Present ages.
Daughter - 12yrs old
MDL - 36yrs old.
Hope this helps.
:-)